# Detecting Out-of-Distribution Objects through Class-Conditioned Inpainting

### v3

@@ 6 Ablation Studies
Denoising process (steps) COCO / OpenImages Avg runtime/image (seconds) FPR@95 ($\downarrow$) AUROC ($\uparrow$) Kandinsky 2.1 20 28.99 / 23.96 89.92 / 91.62 1.12 Stable Diffusion 1 20 29.90 / 23.70 91.36 / 92.36 1.20 Stable Diffusion 2 (our default) 20 25.84 / 18.91 92.28 / 93.30 0.83 Stable Diffusion 2 5 29.48 / 20.20 90.39 / 92.32 0.19 InstaFlow 1 31.34 / 25.74 88.03 / 90.35 0.03
@@
Reliability of Inpainting.   To assess the stability of RONIN independent of specific diffusion inpainting outcomes, we conducted experiments replacing the default Stable Diffusion 2 with Kandinsky 2.1 and Diffusion Model 1, all using 20 denoising steps. We further included InstaFlow (Liu et al., 2024b) with one-step text-to-image generation, and Stable Diffusion 2 with only 5 steps to assert RONIN with low-step performance. As shown in Table 4, RONIN remains stably robust across different diffusion models and settings. Notably, Stable Diffusion 2 offers the best trade-off between performance and runtime, even with as few as 5 steps. These results suggest that, by subtly exploiting the semantic misalignment between original and synthesized objects for zero-shot OOD detection, our proposed RONIN performs consistently, without relying on any specific diffusion model or high-quality inpainted outcomes. Even when diffusion models produce different inpainting results, our framework remains robust and effective.
Table 5: Ablation studies on the triplet similarity measurement (Equation 4). Results show that all three similarities are essential for providing a stable and effective triplet score (underlined). Additionally, prioritizing the vision-language similarity enhances the performance (bolded).
COCO / OpenImages FPR@95 ($\downarrow$) AUROC ($\uparrow$) without $similarity(inp,\hat{y})$ 44.54 / 37.39 89.77 / 91.42 without $similarity(ori,\hat{y})$ 65.15 / 68.91 69.72 / 78.53 without $similarity(ori,inp)$ 30.93 / 36.96 86.14 / 90.10 triplet similarity, $\bm{\alpha}=\bm{\beta}=1$ 35.46 / 24.78 87.75 / 91.43 triplet similarity, $\bm{\alpha}=1$; $\bm{\beta}=2$ 43.93 / 33.91 85.40 / 89.50 triplet similarity, $\bm{\alpha}=2$; $\bm{\beta}=1$ 25.36 / 18.91 92.31 / 93.10 triplet similarity, $\bm{\alpha}=2$; $\bm{\beta}=2$ 35.46 / 25.43 90.32 / 92.32
Triplet Similarity Measurement.    Table 5 presents ablation studies on the components and hyperparameters of the triplet similarity measurement defined in Equation 4. The upper half of the table (rows 1–4) evaluates the impact of removing individual similarity components, highlighting the necessity of incorporating all components to achieve optimal performance. The lower half (rows 5–7) investigates the balance between semantic alignment ($\textit{similarity}{(ori,\ \hat{y})}$, weighted by $\bm{\alpha}$) and visual alignment ($\textit{similarity}{(ori,\ inp)}$, weighted by $\bm{\beta}$). Given the semantic gap between ID and OOD cases illustrated in Figure 3, emphasizing semantic alignment ($\bm{\alpha}>\bm{\beta}$) leads to optimal performance, resulting in our design choice. For simplicity, we adopt $\bm{\alpha}=2$ and $\bm{\beta}=1$ as our default configuration. However, we hypothesize that the sensitive natures of these weights allow RONIN to adapt flexibly to more challenging scenarios.
7 Discussion
Through consistently strong performance across evaluated experiments and ablation studies, our proposed framework, RONIN, effectively demonstrates its flexibility, generality, and practicality across a wide range of settings and scenarios, from indoor scenes to outdoor or in-the-wild environments. By effectively leveraging the subtle misalignment between generative and discriminative outputs, our method performs robustly without relying on specific diffusion models or inpainting outcomes, while remaining versatile and adaptable to various object detection models, including open-world ones, in a training-free manner. However, RONIN is not without limitations. The computational cost of diffusion models restricts RONIN applicability in real-time settings. Nevertheless, many real-world tasks often require offline post-processing, where accuracy and generalizability are prioritized over runtime, in which our method continues to excel. Moreover, ongoing advancements in generative modeling, such as one-step diffusion models InstaFlow (Liu et al., 2024b), offer promising improvements in efficiency, further enhancing RONIN practicality.
8 Conclusion
Overall, RONIN offers a plug-and-play solution for OOD object detection, delivering strong and consistent performance across scenarios. Without requiring access to ID data or retraining effort, RONIN is highly adaptable and compatible, leveraging general-purpose pre-trained models to enhance domain-specific OOD detection both practically and logically. By subtly leveraging the inconsistency between generative and discriminative models with a triplet similarity that considers both semantic and visual alignments, RONIN excels across various OOD detection tasks and seamlessly integrates with most existing object detectors for flexible, dynamic use. With consistently robust experimental results, RONIN demonstrates its flexibility, generality, and practicality, contributing meaningfully to the advancement of zero-shot out-of-distribution object detection.
Appendix A Data processing
Following Du et al. (2022b, a), we used PascalVOC (Everingham et al., 2010) with 20 categories [^1] PascalVOC contains the following ID labels: Person, Car, Bicycle, Boat, Bus, Motorbike, Train, Aeroplane, Chair, Bottle, Dining Table, Potted Plant, TV Monitor, Couch, Bird, Cat, Cow, Dog, Horse, Sheep.
and BDD100k (Yu et al., 2020) with 10 categories [^2] BDD100k contains the following ID labels: Pedestrian, Rider, Car, Truck, Bus, Train, Motorcycle, Bicycle, Traffic Light, Traffic Sign. We utilized two subsets of MS-COCO (Lin et al., 2014) and OpenImages (Kuznetsova et al., 2020) as the OOD sets, where all images containing ID categories are removed to ensure the disjoint-category requirement. Additionally, for some British labels such as “Aeroplane” or “Couch”, we replaced them with the US counterparts such as “Airplane” or “Sofa” to optimize the usage of CLIP visual-textual representations.
Since our method is zero-shot and does not require training, all the experiments are conducted in subsets taken from the testing set of VOC, BDD, COCO, and OpenImages. To ensure intensive, fairness, and effective evaluations, as well as the balance between ID and OOD objects, we sample a subset of 200 images for BDD-100k testing set and 400 images for PascalVOC, MS-COCO, and OpenImages, as BDD-100k often contain a significantly larger amount of objects per image.
Appendix B Implementation Details for Baselines
For the discriminative zero-shot approaches, MCM (Ming et al., 2022), CLIPN (Wang et al., 2023), TAG (Liu & Zach, 2024), OLE (Ding & Pang, 2024) and GL-MCM (Miyai et al., 2025) are originally developed for image-level OOD detection; we adapt it to the object-level setting by applying it over the cropped detections. ODIN (Liang et al., 2018) and Energy score (Liu et al., 2020) are originally developed for non-zero-shot classifier-based settings; we adapt them to the zero-shot object-level setting by using CLIP to classify the detection crops. For alternative generative approaches, we first generate synthetic objects using Stable Diffusion 2, and then extract features with SimCLRv2 (Chen et al., 2020) and fit OOD detectors using standard methods like Mahalanobis (Xiao et al., 2021) and KNN (Sun et al., 2022), for a fair comparison. For SIREN (Du et al., 2022a) and VOS (Du et al., 2022b), we follow their original settings, using the features from the detectors trained with their proposed schemes.
Appendix C Ablation on Inpainting Masking Ratio
To investigate the correlation between the predicted bounding box and the inpainting mask, we conduct an experiment where the sampled mask covers 25% to 100% of the bounding box. The results, presented in Table 6 and visualized in Figure 7, demonstrate that RONIN is robust across different masking ratios but achieves optimal performance when 75% to 90% of the object is covered. As shown in Figure 7, lower masking ratios result in inpaintings that closely resemble the original objects in both ID and OOD cases, preserving the semantic structure of the originals. Conversely, masking the entire object produces inpaintings that deviate significantly from the originals in both cases, introducing ambiguity in OOD detection. By covering 75%-90% of the object, the inpaintings retain the visual appearance and semantic information of ID objects while generating dissimilar representations for OOD cases, enabling RONIN to effectively distinguish ID and OOD samples.
Table 6: Ablation studies on the masking ratio. RONIN maintains consistent and robust performance across different masking ratios, with the best performance attained when 75%-90% of each object is masked.
Covered masking MS-COCO/OpenImages
ratio $\bm{m}$ FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ )
25% 42.47 / 36.74 88.53 / 90.40
50% 34.23 / 23.91 89.39 / 92.45
75% 30.10 / 21.74 91.43 / 93.41
80% 26.80 / 16.74 91.32 / 93.84
90% 25.57 / 19.57 91.53 / 92.46
100% 28.87 / 21.09 90.61 / 92.61
Figure 7: RONIN inpainting performance with different masking ratios. A small masking ratio preserves much of the original object’s semantic structure, resulting in high similarity for both ID and OOD cases. In contrast, a high masking ratio introduces large visual deviations, leading to lower similarity in both cases. With a balanced masking ratio around 75–90%, RONIN effectively retains the visual and semantic integrity of ID objects while inducing significant deviations in OOD cases, achieving a clear and distinct separation between them.

### v4

@@ C.1 Impact of Masking Ratios
Denoising Process (steps) COCO / OpenImages Avg runtime/image (seconds) FPR@95 ($\downarrow$) AUROC ($\uparrow$) Many-step Regime Kandinsky 2.1  [^3] https://huggingface.co/kandinsky-community/kandinsky-2-1-inpaint 20 28.99 / 23.96 89.92 / 91.62 1.12 Stable Diffusion 1.5  [^4] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-inpainting 20 29.90 / 23.70 91.36 / 92.36 1.20 Stable Diffusion 2 (default choice) 20 25.84 / 18.91 92.28 / 93.30 0.83 Stable Diffusion 3  [^5] https://huggingface.co/IrohXu/stable-diffusion-3-inpainting 20 21.09 / 18.89 94.01 / 94.43 2.86 Few-step Regime Stable Diffusion 2 5 27.94 / 20.00 92.35 / 92.32 0.19 InstaFlow 1 31.34 / 25.74 88.03 / 90.35 0.03
C.2 Robust across Generative Models
@@
Extending the ablation study in Tab. 6 to the many-step regime, we further evaluate RONIN by replacing the default Stable Diffusion 2 with Kandinsky 2.1 [razzhigaev2023kandinsky] and Diffusion Model 1, both using 20 denoising steps. We also include results from the few-step regime study, with the one-step diffusion model InstaFlow [liu2023instaflow] and 5-step Stable Diffusion 2 for comparison. As shown in Tab. 9, RONIN remains consistently robust across different diffusion models and configurations. Notably, Stable Diffusion 2 provides the best trade-off between performance and runtime, even in the few-step generative mode. These findings suggest that by subtly leveraging semantic misalignment between original and synthesized objects for zero-shot OOD detection, RONIN performs stably and reliably without relying on any specific diffusion model or high-quality inpainted outputs. Even when models produce diverse inpainting results, our framework remains both robust and effective.